# On Stronger Computational Separations Between Multimodal and Unimodal Machine Learning

### v1

@@ Abstract
In multimodal machine learning, multiple modalities of data (e.g., text and images) are combined to facilitate the learning of a better machine learning model, which remains applicable to a corresponding unimodal task (e.g., text generation). Recently, multimodal machine learning has enjoyed huge empirical success (e.g. GPT-4). Motivated to develop theoretical justification for this empirical success, Lu (NeurIPS ’23, ALT ’24) introduces a theory of multimodal learning, and considers possible separations between theoretical models of multimodal and unimodal learning. In particular, Lu (ALT ’24) shows a computational separation, which is relevant to worst-case instances of the learning task.
In this paper, we give a stronger average-case computational separation, where for “typical” instances of the learning task, unimodal learning is computationally hard, but multimodal learning is easy. We then question how “organic” the average-case separation is. Would it be encountered in practice? To this end, we prove that under natural conditions, any given computational separation between average-case unimodal and multimodal learning tasks implies a corresponding cryptographic key agreement protocol. We suggest to interpret this as evidence that very strong computational advantages of multimodal learning may arise infrequently in practice, since they exist only for the “pathological” case of inherently cryptographic distributions. However, this does not apply to possible (super-polynomial) statistical advantages.
@@ 1 Introduction
Theorem 1.3 also provides a hint for why we do not achieve a separation using the weaker standard LPN assumption, where the noise rate $\theta$ can be taken to be any constant fraction less than one half, and the secret parity is uniformly random. Indeed, the standard LPN assumption is not known to imply any form of KA, unlike the low-noise variant. Constructing KA from the standard LPN assumption is a major open problem in cryptography.
@@ Appendix A Bit Agreement
After proving hardness for $\rho_{\rm uni}\in\{\rho_{\mathcal{X},\mathcal{Z}},\rho_{\mathcal{Y},\mathcal{Z}}\}$ suffices to prove the theorem. ∎

### v2

@@ Abstract
Recently, multimodal machine learning has enjoyed huge empirical success (e.g. GPT-4). Motivated to develop theoretical justification for this empirical success, Lu (NeurIPS ’23, ALT ’24) introduces a theory of multimodal learning, and considers possible separations between theoretical models of multimodal and unimodal learning. In particular, Lu (ALT ’24) shows a computational separation, which is relevant to worst-case instances of the learning task. In this paper, we give a stronger average-case computational separation, where for “typical” instances of the learning task, unimodal learning is computationally hard, but multimodal learning is easy. We then question how “natural” the average-case separation is. Would it be encountered in practice? To this end, we prove that under basic conditions, any given computational separation between average-case unimodal and multimodal learning tasks implies a corresponding cryptographic key agreement protocol. We suggest to interpret this as evidence that very strong computational advantages of multimodal learning may arise infrequently in practice, since they exist only for the “pathological” case of inherently cryptographic distributions. However, this does not apply to possible (super-polynomial) statistical advantages.
@@ 1 Introduction
Theorem 1.3 also provides a hint for why we do not achieve a separation as in Theorem 1.2 by using the weaker standard LPN assumption, where the noise rate $\theta$ can be taken to be any constant fraction less than one half, and the secret parity is uniformly random. Indeed, the standard LPN assumption is not known to imply any form of KA, unlike the low-noise variant. Constructing KA from the standard LPN assumption is a major open problem in the theory of cryptography.
@@ 2.5 Average-case Computational Separations
2.6 Relationship to LUPI
The model for multimodal PAC learning defined by Lu (2023a), which is used in this paper, bears resemblance to the Learning Using Privileged Information (LUPI) paradigm of Vapnik and Vashist (2009). In fact, for the bimodal case, the two models of learning are the same (we omit formal proof, which follows immediately by definition).
Both LUPI and bimodal learning consider triplets of information (rather than the standard of pairs). These triplets are given as input to the learning algorithm at training time, while at test time, only pairs are received. In this way, the two models consider situations where multiple modalities are accessible for training machine learning models, and may be effective for learning problems that act on just a single modality at test time. In the case of LUPI, the second modality is motivated by the presence of a “teacher.” The “teacher” can try to ease the learning process by giving the learner some auxiliary information about the data. In the case of multimodal learning, the additional modalities are motivated by the contemporary success of multimodal perception in AI, where there is an abundance of data but not necessarily a “teacher.”
Since the LUPI paradigm of Vapnik and Vashist (2009) is the same as the bimodal learning model of Lu (2023a), both our main results (Theorem 1.2 and 1.3) apply to the LUPI paradigm. Previous work (Vapnik and Vashist, 2009; Lapin et al., 2014; Vapnik et al., 2015) on understanding the LUPI paradigm (e.g. proving upper bounds, lower bounds, and separations) focused on statistical learning settings such as empirical risk minimization (ERM). To our knowledge, our results are the first to consider the average-case computational power of the LUPI paradigm.
@@ 5 Conclusion
Acknowledgements
This work was supported by the DARPA SIEVE program, Agreement Nos. HR00112020021 and HR00112020023. The author thanks Zhou Lu for several helpful conversations, and Sivan Sabato for pointing out a relationship to the LUPI paradigm.
@@ Appendix A Bit Agreement
After proving hardness for $\rho_{\rm uni}\in\{\rho_{\mathcal{X},\mathcal{Z}},\rho_{\mathcal{Y},\mathcal{Z}}\}$, this suffices to prove the theorem. ∎